# Numerical stability of force-gradient integrators and their Hessian-free variants in lattice QCD simulations

### v1

@@ Abstract
We investigate the numerical stability of force-gradient integrators and their Hessian-free variants within the molecular dynamics step of the Hamiltonian Monte Carlo algorithm in lattice QCD simulations. A linear stability analysis of (Hessian-free) force-gradient integrators is conducted by investigating the harmonic oscillator as a test equation. By performing detailed stability investigations for the entire family of self-adjoint integrators with up to eleven exponentials per time step, we detect promising integrator variants that are providing a good trade-off between accuracy and numerical stability. Simulations for the two-dimensional Schwinger model demonstrate that there are no significant differences in the stability domain of a force-gradient integrator and its Hessian-free counterpart. Furthermore, lattice QCD simulations are conducted to emphasize the significance of numerical stability as a metric for evaluating the computational efficiency of integrators when applied to lattice QCD simulations.
@@
MSC:
[2020] 81V05 , 65P10 , 65L05 , 65L20
journal: Elsevier,  Preprint CERN-TH-2025-093
@@ 1 Introduction
The Hamiltonian Monte Carlo (HMC) algorithm [duane1987hybrid] is a frequently employed algorithm for lattice QCD simulations. In the molecular dynamics (MD) step of the HMC algorithm, trial configurations are generated by integrating the Hamiltonian equations of motion from an initial configuration for a fictitious time $\tau$. In a subsequent Metropolis step [metropolis1953], configurations are accepted or rejected by investigating the energy change $\Delta\mathcal{H}$ along a trajectory.
The selection of the numerical integration scheme is crucial for an efficient and structure-preserving integration of the equations of motion. Firstly, the numerical integration scheme must be volume-preserving and time-reversible to satisfy the detailed balance condition [duane1987hybrid], ensuring that the Markov process converges to the correct equilibrium. Secondly, the links are elements of the special unitary group $\mathrm{SU}(3)$, and therefore we additionally require that the integrator provides numerical approximations situated on the Lie group manifold.
@@ 1.1 Decomposition algorithms
For separable Hamiltonian systems $\mathcal{H}(\mathbf{q},\mathbf{p})=\mathcal{T}(\mathbf{p})+\mathcal{V}(\mathbf{q})$, decomposition algorithms [mclachlan2002splitting, omelyan2003symplectic] facilitate the derivation of explicit numerical integration schemes that adhere to the aforementioned properties. In addition to conventional splitting methods [mclachlan2002splitting], force-gradient integrators [omelyan2003symplectic] enhance the computational process by incorporating a symplectic corrector [wisdom1996symplectic] into the momentum updates. The correction necessitates the inclusion of the force-gradient term, which involves the Hessian of the potential. Hessian-free force-gradient integrators [schäfers2024hessianfree] utilize an approximation of the force-gradient updates, effectively replacing the force-gradient term in the force-gradient step with a second evaluation of the force $-\mathcal{V}_{\mathbf{q}}(\mathbf{q})$. The efficiency of the Hessian-free variants has been demonstrated in various large-scale lattice QCD simulations, see e.g. [Jung2017xef, Finkenrath2022eon, Finkenrath2023sjg, schäfers2024hessianfree].
@@ 1.2 Tuning integrators
When tuning a numerical integration scheme for lattice QCD simulations, we aim at achieving the optimal acceptance rate of $\langle P_{\mathrm{acc}}\rangle_{\mathrm{opt}}=\exp(-1/n)$ with $n$ denoting the convergence order of the numerical integration scheme [TAKAISHI20006]. For normally distributed $\Delta\mathcal{H}$, the acceptance rate corresponds to the variance $\sigma^{2}(\Delta\mathcal{H})$ via
@@
where $\mathrm{erfc}$ denotes the complementary error function [KNECHTLI20033]. Consequently, one aims at maximizing the step size $h$ while satisfying
@@ 1.3 Numerical stability
During the tuning process, the numerical stability of the integrator assumes paramount importance. Decomposition algorithms applied to separable Hamiltonian systems are only conditionally stable, necessitating the selection of a sufficiently small step size $h$. The primary challenge lies in identifying whether the accuracy or the numerical stability of the numerical integration scheme becomes the limiting factor. In such instances, a superior integrator can be obtained by selecting an algorithm that is more accurate or more stable, respectively.
While the accuracy of decomposition algorithms has been extensively studied [omelyan2003symplectic, schäfers2024hessianfree], the numerical stability of decomposition algorithms has only been investigated for conventional splitting methods [BlanesStability]. Since (Hessian-free) force-gradient integrators turn out to be more efficient decomposition algorithms for increasing lattice volumes, an investigation of the numerical stability of (Hessian-free) force-gradient integrators is of particular interest. It is commonly argued [EDWARDS1997375, instabilityMD] that the instability in lattice QCD can be likened to that of a collection of oscillator modes, with the onset of the instability caused by the mode of highest frequency, $\omega_{\max}$. Based on this hypothesis, the linear stability analysis of decomposition algorithms when applied to the harmonic oscillator is sufficient to study the numerical stability for lattice QCD simulations.
@@ 1.4 Outline
The paper is organized as follows. Section 2 provides a brief survey of splitting methods and (Hessian-free) force-gradient integrators. We moreover show that force-gradient integrators and their Hessian-free variants are equivalent when applied to linear systems. Consequently, the linear stability of both frameworks is identical. Subsequently, Section 3 will undertake a linear stability analysis of force-gradient integrators by tailoring the existing linear stability analysis for conventional splitting methods [BlanesStability] to the force-gradient approach. Section LABEL:sec:detection will address the optimization of the algorithms with respect to stability and accuracy. It will be demonstrated that already small alterations to the integrator coefficients will result in a significant reduction in the efficiency measures employed in [omelyan2003symplectic, schäfers2024hessianfree]. Without making any further assumptions on the potential $\mathcal{V}(\mathbf{q})$, the most promising integrators are identified by computing the stability thresholds of all variants listed in [omelyan2003symplectic, schäfers2024hessianfree] that are minimizing the norm of the leading error coefficients. As already mentioned, the linear stability of force-gradient integrators and their Hessian-free variants is identical. In contrast, for nonlinear systems such as those encountered in lattice QCD simulations, the two frameworks differ, potentially leading to differences in numerical stability. Consequently, the numerical tests presented in Section LABEL:sec:numerical_tests start with an examination of potential discrepancies in the numerical stability of force-gradient integrators and their Hessian-free counterparts. This investigation is conducted within the two-dimensional Schwinger model, where an analytical expression of the force-gradient term is available [shcherbakov2017adapted]. These findings substantiate the absence of significant differences in the numerical stability. Afterwards, we will consider an ensemble with two heavy Wilson fermions [Knechtli2022, schäfers2024hessianfree] and demonstrate that the performance of the integrators is consistent with the stability analysis presented in this paper. Furthermore, numerical simulations will be conducted in twisted-mass simulations, providing numerical evidence that the linear stability analysis provides a reliable stability criterion for lattice QCD simulations.
@@ 2 A brief survey of decomposition algorithms
with symmetric and positive definite matrix $\mathbf{M}\in\mathbb{R}^{d\times d}$. Since we will apply the integrators to the harmonic oscillator to perform the linear stability analysis, we consider the canonical case with phase space $\mathbb{R}^{2d}$ here. The formulations of the integrators when the phase space is the cotangent bundle $T^{*}\mathcal{G}$ over a configuration space manifold $\mathcal{G}$ that is a semi-simple matrix Lie group can be found in LABEL:app:LieGroup. For a more general formulation, we refer to [schäfers2024hessianfree, kennedy2013shadow].
@@ 2.1 Hamiltonian equations of motion
$\hat{\mathcal{T}}$ $=\sum_{i,j=1}^{d}m_{i,j}p_{j}\frac{\partial}{\partial q_{i}},$ $\hat{\mathcal{V}}$ $=\sum_{i=1}^{d}\mathcal{V}_{q_{i}}(\mathbf{q})\frac{\partial}{\partial p_{i}}.$ (3)
@@
$\dot{\mathbf{q}}$ $=\hat{\mathcal{T}}\mathbf{q}=\mathbf{M}\mathbf{p},$ $\dot{\mathbf{p}}$ $=-\hat{\mathcal{V}}\mathbf{p}=-\mathcal{V}_{\mathbf{q}}(\mathbf{q}).$
@@ 2.2 Splitting methods
In light of the splitting $\mathcal{H}=\mathcal{T}+\mathcal{V}$, splitting methods [mclachlan2002splitting] compute a numerical approximation of the formal solution $\operatorname{\mathbf{e}}^{{h}(\hat{\mathcal{T}}+\hat{\mathcal{V}})}$ through a composition of the position updates (5) and momentum updates (6). An $s$-stage splitting method is defined by arbitrary coefficients $a_{1},b_{1},\ldots,a_{s},b_{s}\in\mathbb{R}$ and can be expressed as
@@
$\sum_{\ell=1}^{d}a_{\ell}=\sum_{\ell=1}^{d}b_{\ell}=1,$ (9)
the splitting method is consistent and, due to the inherent symmetry, of convergence order $n\geq 2$. The order conditions for splitting methods can be determined by applying the Baker–Campbell–Hausdorff (BCH) formula [mclachlan2002splitting]. A splitting method (7) is of convergence order $n$ if
@@
For further details on the order theory and the construction of higher-order splitting methods, we refer to [mclachlan2002splitting, Blanes_Casas_Murua_2024] and the references therein.
@@ 2.3 Force-gradient integrators
$\begin{split}&\operatorname{\mathbf{e}}^{b_{\ell}{h}\hat{\mathcal{V}}+c_{\ell}{h}^{3}\hat{\mathcal{C}}}\begin{pmatrix}\mathbf{q}_{0}\\ \mathbf{p}_{0}\end{pmatrix}\\ &\qquad\coloneqq\begin{pmatrix}\mathbf{q}_{0}\\ \mathbf{p}_{0}-b_{\ell}{h}\mathcal{V}_{\mathbf{q}}(\mathbf{q}_{0})+2c_{\ell}{h}^{3}\mathcal{V}_{\mathbf{q}\mathbf{q}}(\mathbf{q}_{0})\mathbf{M}\mathcal{V}_{\mathbf{q}}(\mathbf{q}_{0})\end{pmatrix}.\end{split}$ (11)
A force-gradient integrator [kennedy2009force, omelyan2003symplectic] with $s$ stages reads
@@
A force-gradient step (11) involving the force-gradient term $\hat{\mathcal{C}}$ may be regarded as a momentum update (6), applied to a Hamiltonian with modified potential
@@
where $\mathcal{C}(\mathbf{q})$ is solved by $\frac{\partial}{\partial q_{i}}\mathcal{C}(\mathbf{q})=2\sum_{j,k=1}^{d}\mathcal{V}_{q_{k}}(\mathbf{q})m_{j,k}\mathcal{V}_{q_{i}q_{j}}(\mathbf{q})$. Moreover, the modified potential is again Hamiltonian, i.e., the exact flow is again a symplectic map. Thus, force-gradient integrators are symplectic and, as a consequence, volume-preserving. Self-adjoint force-gradient integrators, that is, those integrators satisfying the symmetry relations (8) and where the coefficients $c_{\ell}$ moreover satisfy the same symmetry relations as the coefficients $b_{\ell}$, are time-reversible. For a comprehensive analysis of the order theory of force-gradient integrators, including a detailed classification of self-adjoint integrators with $s\leq 6$ stages, we refer to [omelyan2003symplectic].
@@ 2.4 Hessian-free force-gradient integrators
If computing the force-gradient update (11) is significantly more expensive than computing a momentum update (6), or if an expression for (10) is not available (either analytically or in a software implementation), then Hessian-free force-gradient integrators [schäfers2024hessianfree] represent a promising alternative. In this approach, the force-gradient step (11) is approximated via
@@
effectively replacing the Hessian of the potential by an additional force evaluation. Consequently, an $s$-stage Hessian-free force-gradient integrator is given by
@@
The approximation of the force-gradient step does not affect the time-reversibility of the method, provided that the composition is self-adjoint. As a composition of shears, Hessian-free force-gradient integrators are also volume-preserving. The approximated force-gradient step (13), however, is no longer a symplectic map, except in the special cases where either $c_{\ell}=0$ or the Hessian of the potential commutes with itself when evaluated at (in general, different) $\mathbf{q}$. A consequence of symplecticity is the preservation of a nearby shadow Hamiltonian [kennedy2013shadow]. In contrast to force-gradient integrators, the Hessian-free variants do not preserve a shadow Hamiltonian. In general, without any additional assumptions on the potential, the numerical energy will thus exhibit a drift of size $\mathcal{O}(\tau{h}^{\max\{4,n\}})$. The lack of long-time energy conservation of the Hessian-free variants may prove disadvantageous, particularly for exponentially long-time simulations. In lattice QCD, however, trajectory lengths of $\tau\approx 2$ are typically considered. Therefore, the theoretical energy drift has no impact on the acceptance probability of the HMC algorithm. A Taylor series expansion of the approximation (13) reveals that the additional error terms involve derivatives of the potential $\mathcal{V}$ of order three and higher. It follows that for quadratic Hamiltonian functions, i.e., Hamiltonian functions (2) with potential energy $\mathcal{V}(\mathbf{q})=\frac{1}{2}\mathbf{q}^{\top}\mathbf{N}\mathbf{q}$, the approximated force-gradient update (13) is identical to the force-gradient step (11). As a consequence, the two frameworks are equivalent when applied to linear ODE systems like the harmonic oscillator, i.e., their linear stability is identical.

### v2

@@ Abstract
A comprehensive linear stability analysis of force-gradient integrators and their Hessian-free variants is carried out by investigating the harmonic oscillator as a test equation. The analysis reveals that the linear stability of conventional force-gradient integrators and their Hessian-free counterparts coincides. By performing detailed linear stability investigations for the entire family of self-adjoint integrators with up to eleven exponentials per time step, we detect promising integrator variants that are providing a good trade-off between accuracy and numerical stability. Special attention is given to the application of these promising integrator variants within the Hamiltonian Monte Carlo algorithm, particularly in the context of interacting field theories. Simulations for the two-dimensional Schwinger model are conducted to demonstrate that there are no significant differences in the stability domain of a force-gradient integrator and its Hessian-free counterpart. Lattice QCD simulations with two heavy Wilson fermions emphasize that Hessian-free force-gradient integrators with a larger stability threshold allow for a more efficient computational process compared to conventional splitting methods. Furthermore, detailed investigations of the stability threshold are performed by investigating $N_{f}=2$ twisted-mass fermions and nested integrators, highlighting the reliability of the linear stability threshold for lattice QCD simulations.
@@
2020 MSC:
81V05 , 65P10 , 65L05 , 65L20
journal: Computer Physics Communications
@@ 1 Introduction
The Hamiltonian Monte Carlo (HMC) algorithm duane1987hybrid is a frequently employed algorithm for lattice QCD simulations. In the molecular dynamics (MD) step of the HMC algorithm, trial configurations are generated by integrating the Hamiltonian equations of motion from an initial configuration for a fictitious time $\tau$. In a subsequent Metropolis step metropolis1953 , configurations are accepted or rejected by investigating the energy change $\Delta\mathcal{H}$ along a trajectory.
The selection of the numerical integration scheme is crucial for an efficient and structure-preserving integration of the equations of motion. Firstly, the numerical integration scheme must be volume-preserving and time-reversible to satisfy the detailed balance condition duane1987hybrid , ensuring that the Markov process converges to the correct equilibrium. Secondly, the links are elements of the special unitary group $\mathrm{SU}(3)$, and therefore we additionally require that the integrator provides numerical approximations situated on the Lie group manifold.
@@ 1.1 Decomposition algorithms
For separable Hamiltonian systems $\mathcal{H}(\mathbf{q},\mathbf{p})=\mathcal{T}(\mathbf{p})+\mathcal{V}(\mathbf{q})$, decomposition algorithms mclachlan2002splitting , omelyan2003symplectic facilitate the derivation of explicit numerical integration schemes that adhere to the aforementioned properties. In addition to conventional splitting methods mclachlan2002splitting , force-gradient integrators omelyan2003symplectic enhance the computational process by incorporating a symplectic corrector wisdom1996symplectic into the momentum updates. The correction necessitates the inclusion of the force-gradient term, which involves the Hessian of the potential. Hessian-free force-gradient integrators schäfers2024hessianfree utilize an approximation of the force-gradient updates, effectively replacing the force-gradient term in the force-gradient step with a second evaluation of the force $\mathcal{V}_{\mathbf{q}}(\mathbf{q})$. The efficiency of the Hessian-free variants has been demonstrated in various large-scale lattice QCD simulations, see e.g. Jung2017xef , Finkenrath2022eon , Finkenrath2023sjg , schäfers2024hessianfree .
@@ 1.2 Tuning integrators
When tuning a numerical integration scheme for lattice QCD simulations, we aim at minimizing the computational cost to achieve the optimal acceptance rate of $\langle P_{\mathrm{acc}}\rangle_{\mathrm{opt}}=\exp(-1/n)$ with $n$ denoting the convergence order of the numerical integration scheme TAKAISHI20006 . For normally distributed $\Delta\mathcal{H}$, the acceptance rate corresponds to the variance $\sigma^{2}(\Delta\mathcal{H})$ via
@@
where $\mathrm{erfc}$ denotes the complementary error function KNECHTLI20033 . Consequently, one aims at maximizing the step size h{h} while satisfying
@@ 1.3 Numerical stability
During the tuning process, the numerical stability of the integrator assumes paramount importance. Decomposition algorithms applied to separable Hamiltonian systems are only conditionally stable, necessitating the selection of a sufficiently small step size h{h}. The primary challenge lies in identifying whether the accuracy or the numerical stability of the numerical integration scheme becomes the limiting factor. In such instances, a superior integrator can be obtained by selecting an algorithm that is more accurate or more stable, respectively.
While the accuracy of decomposition algorithms has been extensively studied omelyan2003symplectic , schäfers2024hessianfree , the numerical stability of decomposition algorithms has only been investigated for conventional splitting methods BlanesStability . Since (Hessian-free) force-gradient integrators turn out to be more efficient decomposition algorithms for increasing lattice volumes, an investigation of the numerical stability of (Hessian-free) force-gradient integrators is of particular interest. It is commonly argued EDWARDS1997375 , instabilityMD that the instability in lattice QCD can be likened to that of a collection of oscillator modes, with the onset of the instability caused by the mode of highest frequency, $\omega_{\max}$. Based on this hypothesis, the linear stability analysis of decomposition algorithms when applied to the harmonic oscillator is sufficient to study the numerical stability for lattice QCD simulations.
@@ 1.4 Outline
The paper is organized as follows. Section 2 provides a brief survey of splitting methods and (Hessian-free) force-gradient integrators. We moreover show that force-gradient integrators and their Hessian-free variants are equivalent when applied to linear systems. Consequently, the linear stability of both frameworks is identical. Subsequently, Section 3 will undertake a linear stability analysis of force-gradient integrators by tailoring the existing linear stability analysis for conventional splitting methods BlanesStability to the force-gradient approach. Section 4 will address the optimization of the algorithms with respect to stability and accuracy. It will be demonstrated that already small alterations to the integrator coefficients will result in a significant reduction in the efficiency measures employed in omelyan2003symplectic , schäfers2024hessianfree . Without making any further assumptions on the potential $\mathcal{V}(\mathbf{q})$, the most promising integrators are identified by computing the stability thresholds of all variants listed in omelyan2003symplectic , schäfers2024hessianfree that are minimizing the norm of the leading error coefficients. As already mentioned, the linear stability of force-gradient integrators and their Hessian-free variants is identical. In contrast, for nonlinear systems such as those encountered in lattice QCD simulations, the two frameworks differ, potentially leading to differences in numerical stability. Consequently, the numerical tests presented in Section 5 start with an examination of potential discrepancies in the numerical stability of force-gradient integrators and their Hessian-free counterparts. This investigation is conducted within the two-dimensional Schwinger model, where an analytical expression of the force-gradient term is available shcherbakov2017adapted . These findings substantiate the absence of significant differences in the numerical stability. Afterwards, we will consider an ensemble with two heavy Wilson fermions Knechtli2022 , schäfers2024hessianfree and demonstrate that the performance of the integrators is consistent with the stability analysis presented in this paper. Furthermore, numerical simulations will be conducted with $N_{f}=2$ twisted-mass fermions, providing numerical evidence that the linear stability analysis provides a reliable stability criterion for lattice QCD simulations.
@@ 2 A brief survey of decomposition algorithms
with symmetric and positive definite matrix $\mathbf{M}\in\mathbb{R}^{d\times d}$. Since we will apply the integrators to the harmonic oscillator to perform the linear stability analysis, we consider the canonical case with phase space $\mathbb{R}^{2d}$ here. The formulations of the integrators when the phase space is the cotangent bundle $T^{*}\mathcal{G}$ over a configuration space manifold $\mathcal{G}$ that is a semi-simple matrix Lie group can be found in A. For a more general formulation, we refer to schäfers2024hessianfree , kennedy2013shadow .
@@ 2.1 Hamiltonian equations of motion
$\hat{\mathcal{T}}$ $=\sum_{i,j=1}^{d}m_{i,j}p_{j}\frac{\partial}{\partial q_{i}},$ $\hat{\mathcal{V}}$ $=-\sum_{i=1}^{d}\mathcal{V}_{q_{i}}(\mathbf{q})\frac{\partial}{\partial p_{i}}.$ (3)
@@
$\dot{\mathbf{q}}$ $=\hat{\mathcal{T}}\mathbf{q}=\mathbf{M}\mathbf{p},$ $\dot{\mathbf{p}}$ $=\hat{\mathcal{V}}\mathbf{p}=-\mathcal{V}_{\mathbf{q}}(\mathbf{q}).$
@@ 2.2 Splitting methods
In light of the splitting $\mathcal{H}=\mathcal{T}+\mathcal{V}$, splitting methods mclachlan2002splitting compute a numerical approximation of the formal solution $\operatorname{\mathbf{e}}^{{h}(\hat{\mathcal{T}}+\hat{\mathcal{V}})}$ through a composition of the position updates (5) and momentum updates (6). A $s$-stage splitting method is defined by arbitrary coefficients $a_{1},b_{1},\ldots,a_{s},b_{s}\in\mathbb{R}$ and can be expressed as
@@
$\sum_{\ell=1}^{s}a_{\ell}=\sum_{\ell=1}^{s}b_{\ell}=1,$ (9)
the splitting method is consistent and, due to the inherent symmetry, of convergence order $n\geq 2$. The order conditions for splitting methods can be determined by applying the Baker–Campbell–Hausdorff (BCH) formula mclachlan2002splitting . A splitting method (7) is of convergence order $n$ if
@@
For further details on the order theory and the construction of higher-order splitting methods, we refer to mclachlan2002splitting , Blanes_Casas_Murua_2024 and the references therein.
@@ 2.3 Force-gradient integrators
[…]
A force-gradient integrator kennedy2009force , omelyan2003symplectic with $s$ stages reads
@@
A force-gradient step (11) involving the force-gradient term $\hat{\mathcal{C}}$ may be regarded as a conventional momentum update (6), applied to a Hamiltonian with modified potential
@@
where $\mathcal{C}(\mathbf{q})$ is solved by $\frac{\partial}{\partial q_{i}}\mathcal{C}(\mathbf{q})=2\sum_{j,k=1}^{d}\mathcal{V}_{q_{k}}(\mathbf{q})m_{j,k}\mathcal{V}_{q_{i}q_{j}}(\mathbf{q})$. Moreover, the modified potential is again Hamiltonian, i.e., the exact flow is again a symplectic map. Thus, force-gradient integrators are symplectic and, as a consequence, volume-preserving. Self-adjoint force-gradient integrators, that is, those integrators satisfying the symmetry relations (8) and where the coefficients $c_{\ell}$ moreover satisfy the same symmetry relations as the coefficients $b_{\ell}$, are time-reversible. For a comprehensive analysis of the order theory of force-gradient integrators, including a detailed classification of self-adjoint integrators with $s\leq 6$ stages, we refer to omelyan2003symplectic .
@@ 2.4 Hessian-free force-gradient integrators
If computing the force-gradient update (11) is significantly more expensive than computing a momentum update (6), or if an expression for (10) is not available (either analytically or in a software implementation), then Hessian-free force-gradient integrators schäfers2024hessianfree represent a promising alternative. In this approach, the force-gradient step (11) is approximated via
@@
effectively replacing the Hessian of the potential by an additional force evaluation. Consequently, a $s$-stage Hessian-free force-gradient integrator is given by
@@
The approximation of the force-gradient step does not affect the time-reversibility of the method, provided that the composition is self-adjoint. As a composition of shears, Hessian-free force-gradient integrators are also volume-preserving. The approximated force-gradient step (13), however, is no longer a symplectic map, except in the special cases where either $c_{\ell}=0$ or the Hessian of the potential commutes with itself when evaluated at (in general, different) $\mathbf{q}$. A consequence of symplecticity is the preservation of a nearby shadow Hamiltonian kennedy2013shadow . In contrast to force-gradient integrators, the Hessian-free variants do not preserve a shadow Hamiltonian. In general, without any additional assumptions on the potential, the numerical energy will thus exhibit a drift of size $\mathcal{O}(\tau{h}^{\max\{4,n\}})$. The lack of long-time energy conservation of the Hessian-free variants may prove disadvantageous, particularly for exponentially long-time simulations. In lattice QCD, however, trajectory lengths of $\tau\approx 2$ are typically considered. Therefore, the theoretical energy drift has no impact on the acceptance probability of the HMC algorithm. A Taylor series expansion of the approximation (13) reveals that the additional error terms involve derivatives of the potential $\mathcal{V}$ of order three and higher. It follows that for quadratic Hamiltonian functions, i.e., Hamiltonian functions (2) with potential energy $\mathcal{V}(\mathbf{q})=\frac{1}{2}\mathbf{q}^{\top}\mathbf{N}\mathbf{q}$, the approximated force-gradient update (13) is identical to the force-gradient step (11). As a consequence, the two frameworks are equivalent when applied to linear ODE systems like the harmonic oscillator, i.e., their linear stability is identical.